# MiMiC: A High-Performance Framework for Multiscale Molecular Dynamics Simulations

### v1

@@ Abstract
MiMiC is a framework for performing multiscale simulations, where individual subsystems are handled at different resolutions and/or levels of theory by loosely coupled external programs. To make it highly efficient and flexible, we adopt an interoperable approach based on a multiple-program multiple-data paradigm, serving as an intermediary responsible for fast data exchange and interactions between the subsystems. The main goal of MiMiC is to avoid interfering with the underlying parallelization of the external programs, including the operability on hybrid architectures (e.g., CPU/GPU), and keep their setup and execution as close as possible to the original. At the moment, MiMiC offers an efficient implementation of electrostatic embedding QM/MM that has demonstrated unprecedented parallel scaling in simulations of large biomolecules using CPMD and GROMACS as QM and MM engines, respectively. However, as it is designed for high flexibility with general multiscale models in mind, it can be straightforwardly extended beyond QM/MM. In this article, we illustrate the software design and the features of the framework, which make it a compelling choice for multiscale simulations in the upcoming era of exascale high-performance computing.
@@ I Introduction
Multiscale methods are usually implemented by extending a program with the functionality of another, e.g., in the case of QM/MM, a QM program with MM functionality, or vice versa. These implementations range from fully monolithic, through linked libraries with ad hoc interfaces to stand-alone programs, up to general interfaces between independent programs. Although most QM/MM implementations adopt the first two of these strategies, some software packages implement more flexible interfaces that allow coupling to, in principle, any other program without substantial effort.Götz, Clark, and Walker (2014); Melo et al. (2018); Cruzeiro et al. (2023) Apart from these already more general interfaces, an even more flexible approach is to implement multiscale methods through integrative frameworks. These do not themselves implement any QM, MM, or other such functionality for calculating the properties of individual subsystems, but instead rely completely on external programs.Woodcock et al. (2011); Ma et al. (2015); Torras, Deumens, and Trickey (2006); Torras et al. (2015); Řezáč (2016); Larsen et al. (2017); Altoè et al. (2007); Weingart et al. (2018); Metz et al. (2014); Lu et al. (2019); Kratz et al. (2016); Gökcan, Vázquez-Montelongo, and Cisneros (2019); Zhang et al. (2019); Barnes et al. (2021); Martí (2021); Lin et al. (2024)
@@
An important aspect of multiscale implementations involving several distinct software components is their coupling, which can be classified as tight or loose based on the degree of interdependency between them. Loose coupling helps to avoid code duplication and greatly reduces maintenance costs, as modifications made to one component do not impact the rest and conversely, new features implemented in a given component are instantly available with practically no additional effort. In addition, replacing individual components becomes almost trivial. Loose coupling to the external programs is fundamental to essentially all general multiscale frameworks. However, obtaining high computational efficiency can pose a challenge due to potentially slow data exchange compared to the tight-coupling approach, where the relevant data is available in memory. In particular, most frameworks rely on file-based communication, i.e., using input and output files, which can have a substantial impact on performance.Isborn et al. (2012); Cruzeiro et al. (2021) To overcome these issues, some adopt a library-linking approach that requires converting the external programs into libraries, thus losing some of the flexibility, while others are realized through network-based communication.
@@
Recently, we introduced MiMiCOlsen et al. (2019), a high-performance multiscale modeling framework that combines the best of both worlds, the flexibility of loose coupling, while achieving high computational efficiencyBolnykh et al. (2019). It attains these advantages by employing a client-server approach together with a multiple-program multiple-data (MPMD) model through which it loosely couples external programs that concurrently calculate different subsystem contributions while MiMiC itself calculates interactions between them, see illustration in Figure 1. MiMiC has a dual-sided application programming interface (API), one side of which is used to communicate with a molecular dynamics (MD) driver that integrates the equations of motion, while the other side interfaces to external programs. To ensure efficient and straightforward interfacing with the latter, we developed the MiMiC communication library (MCL) that features a simple API and, moreover, provides a means to achieve efficient network-based communication while minimizing the impact on the underlying parallelization of the external programs. This allows us to exploit the efficiency and features of specialized software packages, thus granting access to multiscale simulations that can take full advantage of state-of-the-art high-performance computing (HPC) architectures.
@@
In this article, we give a detailed presentation of the software design of the MiMiC framework and showcase its features, performance, and recent applications of large biomolecular systems. Section II describes the internal organization of the code and the key aspects of communication, which might also prove useful in potential future interfaces. Section III illustrates workflows of multiscale simulations with MiMiC, focusing on electrostatic and polarizable embedding QM/MM schemes and a multiple time step (MTS) approach for QM/MM MD. Section IV provides an overview of current and upcoming features either implemented directly in MiMiC or accessible through external programs. Section V summarizes some notable studies in which MiMiC enabled the efficient and accurate modeling of complex biophysical and biochemical processes. We conclude this paper by outlining the road ahead.
II Software Design
The design of the MiMiC framework is guided by our primary objectives: achieving optimal flexibility and performance while minimizing implementation and maintenance efforts. Flexibility is attained through a modular design where individual subsystems are assigned to external client programs that compute the relevant properties needed for a multiscale simulation. In this MPMD setup, MiMiC serves as the intermediary that collects and distributes data among the concurrently running client programs, and at the same time, efficiently calculates contributions from the interactions between the subsystems. To accommodate different types of subsystems, particles, and other quantities, we draw extensively on the object-oriented capabilities of modern Fortran and organize data structures into a hierarchy of derived types that can be easily extended. Once a given type of subsystem interaction is implemented, it is then trivially transferable to any compatible program with no additional coding effort besides the implementation of the interface itself. Apart from easy access to new features in the client programs and the MD driver, this approach avoids code duplication and substantially reduces the maintenance effort to keep the MiMiC ecosystem operational.
@@ II.1 Interfaces
An essential feature of the loose-coupling paradigm is the design of a well-defined and general interface between individual programs, as opposed to several program-specific interfaces. For practical purposes, the MiMiC framework is split into two separate libraries, namely, the main MiMiC library and MCL, each providing an API intended for a specific function. The first is linked to the simulation driver program, which also acts as the server that manages communication between the individual clients, keeps track of all simulation-related information, and computes interactions between the subsystems. However, in order to become a client, an external program responsible for a particular subsystem requires only MCL, which facilitates data exchange with the server and can be introduced with only minimal changes in the original code.
@@ II.1.1 Client-Server Communication
To go along with our minimal interference approach for clients, MCL’s API consists of only a handful of procedures, quite similar to those of MPI. Here, we illustrate client-server communication of a generic run within the context of MPI MPMD, see Figure 2.
The simulation starts by launching all programs as an MPI MPMD job, meaning that MPI launches these programs simultaneously. All of them now share a single MPI_COMM_WORLD communicator that is intercepted by invoking MCL_Init() right after the MPI initialization with MPI_Init(). This splits the communicator and returns new local "world" communicators, specific for the individual programs, with all communication between the clients and the server being channeled through intercommunicators. The next step is to establish the client-server connections by calling the MCL_Handshake() procedure in all running programs to specify the executable that will be the server and, at the same time, to identify individual clients by assigning them IDs.
@@
To ensure minimal interference with the infrastructure of a client program, while at the same time maintaining maximum flexibility, we coordinate actions between the clients and the server with a request-based approach. In general, a client initializes as it normally would and then enters a construct that we refer to as the MiMiC loop, which is merely a conditional loop in which two actions are performed every iteration. First, it receives a request via MCL_Recv(), and then it executes the specified action given that it has already been implemented. For an implementation example, see Algorithm 1.
The entire MiMiC loop basically reduces to the execution of one of four possible types of instructions. Requests prepended by MCL_SEND instruct a client to send some sort of data. This could involve allocating a buffer, filling it with requested data, and then sending it using MCL_Send(). The reciprocal action is a receive data request, which is prepended by MCL_RECV. The last type of practical request, denoted by the prefix MCL_RUN, instructs a client to execute a calculation, e.g., to compute the energy, forces, or both. Finally, the server can order clients to stop the program execution by issuing the MCL_EXIT request. Once it is received, the client program wraps up the run, deallocates memory, and ultimately terminates the program. This typically occurs at the end of a simulation or when an error is raised by one of the programs.
@@ II.1.2 MD Driver API
So far, we have described only the client-side API, where MCL is used as a communication channel. However, a program that acts as a server in terms of communication and as a driver in terms of an MD simulation is also needed. In comparison to the client-side interface, this program has to be linked against the MiMiC library, which then provides access to the MiMiC infrastructure via a simple API. It consists of a collection of setter-, getter-, and runner-type procedures that can be easily incorporated into a typical MD loop. At the moment, only a Fortran API is available, but we intend to extend it to other languages.
Let us demonstrate this with a short example of the procedure calls during a single time step of an MD simulation. Once the MD driver updates the particle positions, they are passed to MiMiC using the set_coordinates() function. Then, run_calculation() is called at the point in the code where the MD driver would typically perform a force calculation. By invoking this function, MiMiC carries out all the necessary steps to provide the MD driver with the forces on all particles, i.e., it distributes coordinates to individual clients via MCL, issues commands to start energy and force calculations, collects relevant quantities, and computes interactions between the subsystems. Finally, forces are collected and updated in the MD driver via get_forces(), and the simulation then proceeds with the rest of the MD time step.
@@ II.2 Data Structures
The MiMiC main module comprises two parts, a public API for MD drivers and a private instance of the system class (which here is a Fortran module with a derived type) that contains all the information about the simulation, such as the simulation box parameters, species objects with particle properties, and, most importantly, subsystem and interaction objects. All system data are thus stored in a hierarchy of objects illustrated in Figure 3. This data structure ladder plays a crucial role in the modularity of MiMiC, particularly subsystem and interaction classes, which are the focus of this section.
@@
The interaction class groups objects required for the calculation of subsystem interactions. Here, we use an electrostatic embedding QM/MM scheme as an example to explain the essence of this class. Among its attributes, we include pointers to the involved subsystems, two arrays referencing the short-range (sr) and long-range (lr) MM atoms, and settings related to the atom sorting into sr/lr subdomains (for the theoretical basis, see Section III.1). Finally, its methods constitute an interface to compute different quantities that represent the interaction, be it the potential of MM atoms acting on the QM subsystem, QM/MM energy, or QM/MM forces.
@@ II.3 Parallelization Strategy
MiMiC computes various quantities related to the interactions between subsystems, and it is therefore assigned its own computational resources. To avoid introducing a bottleneck, we use parallelization strategies that are flexible enough to run efficiently on diverse hardware architectures, while at the same time being highly scalable. Therefore, for computationally intensive routines, most data is stored in flat, contiguous arrays, ensuring that the number of nested loops is kept to a minimum. This allows to potentially tailor parallelization to specific needs that might emerge in the future, for example, due to the addition of new subsystem and interaction types.
In terms of employed parallelization schemes, MiMiC currently features a hybrid MPI-based distributed-memory and OpenMP-based shared-memory approach, which is particularly important in the calculation of electrostatic QM/MM interactions, as described in Section III. Within this scheme, the most computationally intensive operation is the numerical integration over the grid storing either an electron density or an external potential, which has to be performed for each MM atom interacting with the QM subsystem. In a typical application, the number of atoms $N_{\mathrm{atom}}$ is on the order of $10^{4}$ to $10^{5}$, while the number of grid points $N_{\mathrm{grid}}$ is on the order of $10^{6}$ to $10^{7}$. However, the number of atoms treated explicitly can be reduced to the order of $10^{3}$ by employing the sr/lr scheme (see Section III.1), thus resulting in a total operation count proportional to $N_{\mathrm{atom}}\times N_{\mathrm{grid}}$, i.e., between $10^{9}$ and $10^{10}$. In the current implementation, it then reduces to two nested loops: the outer loop over the MM atoms is distributed over MPI processes and both loops are further parallelized employing OpenMP-based shared-memory approaches.
@@ II.4 Preprocessing Toolkit
The loose-coupling, minimally-invasive strategy adopted by MiMiC requires the preparation of inputs for multiple external programs that are run concurrently. Sometimes, setting up such simulations can be challenging, especially in view of the partitioning of complex biomolecular systems into two or more subsystems that need to be set up independently. For this purpose, we provide MiMiCPyRaghavan et al. (2023a), a Python-based modular preprocessing toolkit for MiMiC, that supports various formats of topology, coordinates, and input files and is easily extendable to new ones.
@@ II.5 Software Development Practices
The MiMiC framework is free and open-source software distributed under the GNU Lesser General Public License (LGPL) version 3.0 or later. The license was chosen to support the free and open-source software movement without imposing severe restrictions on the client programs since proprietary software is still common in the molecular simulation community. Thus, MCL may be used or integrated into essentially any software without being required to release its source code. The MiMiC source code is hosted on GitLabmim (2024) with releases deposited on ZenodoOlsen et al. (2022); Bolnykh et al. (2023) using versioning according to the semantic versioning specificationPreston-Werner (2013). We strive to maintain a high code quality through code review and continuous integration testing that includes unit and integration tests.

### v2

@@ Abstract
MiMiC is a framework for performing multiscale simulations in which loosely coupled external programs describe individual subsystems at different resolutions and levels of theory. To make it highly efficient and flexible, we adopt an interoperable approach based on a multiple-program multiple-data (MPMD) paradigm, serving as an intermediary responsible for fast data exchange and interactions between the subsystems. The main goal of MiMiC is to avoid interfering with the underlying parallelization of the external programs, including the operability on hybrid architectures (e.g., CPU/GPU), and keep their setup and execution as close as possible to the original. At the moment, MiMiC offers an efficient implementation of electrostatic embedding QM/MM that has demonstrated unprecedented parallel scaling in simulations of large biomolecules using CPMD and GROMACS as QM and MM engines, respectively. However, as it is designed for high flexibility with general multiscale models in mind, it can be straightforwardly extended beyond QM/MM. In this article, we illustrate the software design and the features of the framework, which make it a compelling choice for multiscale simulations in the upcoming era of exascale high-performance computing.
@@ I Introduction
Multiscale methods are usually implemented by extending a program with the functionality of another, e.g., in the case of QM/MM, a QM program with MM functionality, or vice versa. These implementations range from fully monolithic, through linked libraries with ad hoc interfaces to stand-alone programs, up to general interfaces between independent programs. Although most QM/MM implementations adopt the first two of these strategies, some software packages implement more flexible interfaces that allow coupling to, in principle, any other program without substantial effort.Götz, Clark, and Walker (2014); Melo et al. (2018); Cruzeiro et al. (2023) Apart from these already more general interfaces, an even more flexible approach is to implement multiscale methods through integrative frameworks. These do not themselves implement any QM, MM, or other such functionality for calculating the properties of individual subsystems but instead rely completely on external programs.Woodcock et al. (2011); Ma et al. (2015); Torras, Deumens, and Trickey (2006); Torras et al. (2015); Řezáč (2016); Larsen et al. (2017); Altoè et al. (2007); Weingart et al. (2018); Kratz et al. (2016); Gökcan, Vázquez-Montelongo, and Cisneros (2019); Zhang et al. (2019); Barnes et al. (2021); Martí (2021); Metz et al. (2014); Lu et al. (2019, 2023); Lin et al. (2024)
@@
An important aspect of multiscale implementations involving several distinct software components is their coupling, which can be classified as tight or loose based on the degree of interdependency between them. Loose coupling helps to avoid code duplication and greatly reduces maintenance costs, as modifications made to one component do not impact the rest, and replacing individual components becomes almost trivial. Furthermore, new features implemented in a given component are instantly available with practically no additional effort. Loose coupling to the external programs is fundamental to essentially all general multiscale frameworks. However, obtaining high computational efficiency can pose a challenge due to potentially slow data exchange compared to the tight-coupling approach, where the relevant data is available in memory. In particular, most frameworks rely on file-based communication, i.e., using input and output files, which can have a substantial impact on performance.Isborn et al. (2012); Cruzeiro et al. (2021) To overcome these issues, some adopt a library-linking approach that requires converting the external programs into libraries, thus losing some of the flexibility, while others are realized through network-based communication.
@@
Recently, we introduced MiMiCOlsen et al. (2019), a high-performance multiscale modeling framework that combines the best of both worlds, the flexibility of loose coupling, while achieving high computational efficiencyBolnykh et al. (2019). It attains these advantages by employing a client-server approach together with a multiple-program multiple-data (MPMD) model through which it loosely couples external programs that concurrently calculate different subsystem contributions while MiMiC itself calculates interactions between them, see illustration in Fig. 1. MiMiC has a dual-sided application programming interface (API). One side communicates with a molecular dynamics (MD) driver that integrates the equations of motion, while the other interfaces to external programs. To ensure efficient and straightforward interfacing with the latter, we developed the MiMiC communication library (MCL) that features a simple API and, moreover, provides a means to achieve efficient network-based communication while minimizing the impact on the underlying parallelization of the external programs. This allows us to exploit the efficiency and features of specialized software packages, thus granting access to multiscale simulations that can take full advantage of state-of-the-art high-performance computing (HPC) architectures.
In this article, we give a detailed presentation of the software design of the MiMiC framework and showcase its features, performance, and recent applications involving large biomolecular systems. Section II describes the internal organization of the code and the key aspects of communication, which might also prove useful in potential future interfaces. Section III illustrates workflows of multiscale simulations with MiMiC, focusing on electrostatic and polarizable embedding QM/MM schemes and a multiple time step (MTS) approach for QM/MM MD. Sections IV and V provide an overview of current and upcoming features either implemented directly in MiMiC or accessible through external programs. Section VI summarizes some notable studies in which MiMiC enabled the efficient and accurate modeling of complex biophysical and biochemical processes. We conclude this paper by outlining the road ahead.
II Software Design
The design of the MiMiC framework is guided by our primary objectives: achieving optimal flexibility and performance while minimizing implementation and maintenance efforts. Flexibility is attained through a modular design where individual subsystems are assigned to external client programs that compute the relevant properties needed for a multiscale simulation. In this MPMD setup, MiMiC serves as the intermediary that collects and distributes data among the concurrently running client programs and efficiently calculates contributions from the interactions between the subsystems. To accommodate different types of subsystems, particles, and other quantities, we draw extensively on the object-oriented capabilities of modern Fortran and organize data structures into a hierarchy of derived types that can be easily extended. Once a given type of subsystem interaction is implemented, it is then trivially transferable to any compatible program with no additional coding effort besides the implementation of the interface itself. Apart from easy access to new features in the client programs and the MD driver, this approach avoids code duplication and substantially reduces the maintenance effort to keep the MiMiC ecosystem operational.
@@ II.1 Interfaces
An essential feature of the loose-coupling paradigm is the design of a well-defined and general interface between individual programs, as opposed to several program-specific interfaces. For practical purposes, the MiMiC framework is split into two separate libraries, namely, the main MiMiC library and MCL. The first is linked to the simulation driver program, which also acts as the server that manages communication between the individual clients, keeps track of all simulation-related information, and computes subsystem interactions. However, in order to become a client, an external program responsible for a particular subsystem requires only MCL, which facilitates data exchange with the server and can be introduced with only minimal changes in the original code.
@@ II.1.1 Client-Server Communication
To support our minimal interference approach for clients, the MCL API consists of only a handful of procedures similar to those of MPI. Here, we describe client-server communication of a generic run within the context of MPI MPMD, illustrated in Fig. 2.
The simulation starts by launching all programs as an MPI MPMD job, meaning that MPI launches these programs simultaneously. All of them now share a single MPI_COMM_WORLD communicator that is intercepted by invoking MCL_Init() right after the MPI initialization with MPI_Init(). This splits the communicator and returns a new local "world" communicator for each individual program. All communication between the clients and the server is then channeled through intercommunicators. The next step is to establish the client-server connections by calling the MCL_Handshake() procedure in all running programs to specify the executable that will be the server and, at the same time, to identify individual clients by assigning them IDs.
To enable minimal interference with the infrastructure of client programs while maintaining maximum flexibility, we coordinate actions between the clients and the server using a request-based approach. In general, a client initializes as it normally would and then enters a construct that we refer to as the MiMiC loop, which is merely a conditional loop in which two actions are performed in every iteration. First, it receives a request via MCL_Recv(), and then it executes the specified action given that it has already been implemented. For an implementation example, see Algorithm 1.
The entire MiMiC loop basically reduces to the execution of one of four possible types of instructions. Requests prepended by MCL_SEND instruct a client to send the requested data. This could involve allocating a buffer, filling it with requested data, and then sending it using MCL_Send(). The reciprocal action is a receive data request, which is prepended by MCL_RECV. The last type of practical request, denoted by the prefix MCL_RUN, instructs a client to execute a calculation, e.g., to compute the energy, forces, or both. Finally, the server can order clients to stop the program execution by issuing the MCL_EXIT request. Once received, the client program wraps up the run, deallocates memory, and ultimately terminates the program. This typically occurs at the end of a simulation or when one of the programs raises an error.
@@ II.1.2 MD Driver API
So far, we have described only the client-side API, where MCL is used as a communication channel. However, a program that acts as a server in terms of communication and as a driver in terms of an MD simulation is also needed. Unlike the client-side interface, this program has to be linked against the MiMiC library, which then provides access to the MiMiC infrastructure via a simple API. It consists of a collection of setter-, getter-, and runner-type procedures that can be easily incorporated into a typical MD loop. At the moment, only a Fortran API is available, but we intend to extend it to other languages.
Let us demonstrate this with a short example of the procedure calls during a single step of an MD simulation. Once the MD driver updates the particle positions, they are passed to MiMiC using the set_coordinates() function. Then, run_calculation() is called at the point in the code where the MD driver typically performs a force calculation. By invoking this function, MiMiC carries out all the necessary steps to provide the MD driver with the forces on all particles, i.e., it distributes coordinates to individual clients via MCL, issues commands to start energy and force calculations, collects relevant quantities, and computes interactions between the subsystems. Finally, forces are collected and updated in the MD driver via get_forces(), and the simulation proceeds with the rest of the MD step.
@@ II.2 Data Structures
The MiMiC main module comprises two parts, a public API for MD drivers and a private instance of the system class (which here is a Fortran module with a derived type) that contains all the information about the simulation, such as the simulation box parameters, species objects with particle properties, and, most importantly, subsystem and interaction objects. All system data are thus stored in a hierarchy of objects illustrated in Fig. 3. This data structure ladder plays a crucial role in the modularity of MiMiC, particularly subsystem and interaction classes, which are the focus of this section.
@@
The interaction class groups objects required for the calculation of subsystem interactions. Here, we use an electrostatic embedding QM/MM scheme as an example to explain the essence of this class. Among its attributes, we include pointers to the involved subsystems, two arrays referencing the short-range (sr) and long-range (lr) MM atoms, and settings related to the atom sorting into sr/lr subdomains (for the theoretical basis, see Sec. III.1). Finally, its methods constitute an interface for computing different quantities representing the interaction, such as the potential of MM atoms acting on the QM subsystem, QM/MM energy, or QM/MM forces.
@@ II.3 Parallelization Strategy
MiMiC computes various quantities related to the interactions between subsystems, and it is, therefore, assigned its own computational resources. To avoid introducing a bottleneck, we use parallelization strategies that are flexible enough to run efficiently on diverse hardware architectures, while at the same time being highly scalable. Therefore, for computationally intensive routines, most data is stored in flat, contiguous arrays, ensuring that the number of nested loops is kept to a minimum. This allows to potentially tailor parallelization to specific needs that might emerge in the future, such as the addition of new subsystems and interaction types.
In terms of employed parallelization schemes, MiMiC currently features a hybrid MPI-based distributed-memory and OpenMP-based shared-memory approach, which is particularly important in the calculation of electrostatic QM/MM interactions, as described in Sec. III. Within this scheme, the most computationally intensive operation is the numerical integration over the grid storing either an electron density or an external potential, which has to be performed for each MM atom interacting with the QM subsystem. In a typical application, the number of atoms $N_{\mathrm{atom}}$ is on the order of $10^{4}$ to $10^{5}$, while the number of grid points $N_{\mathrm{grid}}$ is on the order of $10^{6}$ to $10^{7}$. However, the number of atoms treated explicitly can be reduced to the order of $10^{3}$ by employing the sr/lr scheme (see Sec. III.1), thus resulting in a total operation count proportional to $N_{\mathrm{atom}}\times N_{\mathrm{grid}}$, i.e., between $10^{9}$ and $10^{10}$. In the current implementation, it then reduces to two nested loops: the outer loop over the MM atoms is distributed over MPI processes and both loops are further parallelized employing OpenMP-based shared-memory approaches.
@@ II.4 Preprocessing Toolkit
The loose-coupling, minimally invasive strategy adopted by MiMiC requires preparing inputs for multiple external programs that are run concurrently. Setting up such simulations can sometimes be challenging, especially since complex biomolecular systems are partitioned into two or more subsystems that need to be set up independently. For this purpose, we provide MiMiCPyRaghavan et al. (2023a), a Python-based modular preprocessing toolkit for MiMiC that supports various formats of topology, coordinates, and input files and is easily extendable to new ones.
@@ II.5 Software Development Practices
The MiMiC framework is free and open-source software distributed under the GNU Lesser General Public License (LGPL) version 3.0 or later. The license was chosen to support the free and open-source software movement without imposing severe restrictions on the client programs since proprietary software is still common in the molecular simulation community. Thus, MCL may be used or integrated into essentially any software without being required to release its source code. The MiMiC source code is hosted on GitLabmim (2024) with releases deposited on ZenodoOlsen et al. (2022); Bolnykh et al. (2023) using versioning according to the semantic versioning specificationPreston-Werner (2013). We strive to maintain high code quality through code review and continuous integration (CI) pipelines, including unit and integration tests.
@@
Figure 4: Demonstration of the efficiency and tunability of MiMiC’s electrostatic embedding scheme for a system consisting of an $n$-butanol molecule (QM subsystem) solvated in acetone (MM subsystem), with a total of 139 495 atoms. The plots show the convergence of the total energy (left) and the forces (center), as well as the relative wall times for the calculation of the QM/MM contributions (right) with respect to the short-range cutoff distance and the order of the multipole expansion (line colors). The points correspond to mean absolute errors in energies, mean of the maximum absolute errors in forces, and mean relative wall times based on single-point calculations on ten snapshots using the full short-range coupling as reference. Data taken from Ref. 34.